Area of a painting. A rectangular painting with a width of x centimeters has an area of x^2 +50x square centimeters. Find a binomial that represents the length.
Answer by richwmiller(17219)

Let y=length
x*y=x^2+50x
x*y=x*(x+50)
y=50+x
Length equals 50+x